OpenStudy (anonymous):

Please help me with this "Proving Identities" 2/1+cos teta + 2/1-cos teta = 4 csc^2 teta

hero (hero):

\[\frac{2}{1 + \cos(\theta)} + \frac{2}{1 - \cos(\theta)} = 4\csc^2(\theta)\]

hero (hero):

^That's what it is, right?

Parth (parthkohli):

@CHAsio_P Any guesses?

Parth (parthkohli):

@Hero Yes.

OpenStudy (anonymous):

@Hero yup @ParthKohli none. how bout you?

hero (hero):

Hint: Combine fractions on the left side.

Parth (parthkohli):

Do you know how to get a common denominator?

OpenStudy (anonymous):

@Hero thanks @ParthKohli well, yeah. So it would be 1- cos \[\Theta r\]

Parth (parthkohli):

You multiply through the denominators, and then you multiply the same to the numerator.\[\dfrac{2\color{#C00}{(1-\cos\theta)}}{\color{#C00}{(1 - \cos\theta)}(1 + \cos\theta)} + \dfrac{2\color{#C00}{(1 + \cos\theta)}}{(1 - \cos\theta)\color{#C00}{(1 + \cos\theta)}}\]

OpenStudy (anonymous):

@ParthKohli how come that the denominator became like that? it is just so confusing. . .

Parth (parthkohli):

Do you know that if you multiply both numerator and denominator by the same number (just not \(0\)), then you get an equivalent fraction.

OpenStudy (anonymous):

@ParthKohli oh, thanks. then, i'll start doing the canceling?

Parth (parthkohli):

No, you will add both fractions.

Parth (parthkohli):

Before that, can you simplify \((1 - \cos\theta)(1 + \cos\theta)\)?

OpenStudy (anonymous):

Honestly, no. Heheh. . .

Parth (parthkohli):

Do you know that \((a - b)(a + b) = a^2 - b^2\)?

OpenStudy (anonymous):

So, the denominator will be \[(1 - \cos ^2 \theta) + (1-\cos^2\theta)\]

OpenStudy (anonymous):

?

Parth (parthkohli):

Yes, you have\[\dfrac{2}{1 -\cos^2 \theta} + \dfrac{2}{1 - \cos^2 \theta}\]

Parth (parthkohli):

When the denominators are equal, you can add the numerators. Consider this elementary example:\[\dfrac{2}{3} + \dfrac{3}{3} = \dfrac{2 + 3}{3} = \dfrac{5}{3}\]

OpenStudy (anonymous):

So, it would be \[\frac{ 2 }{ 1-\cos ^2 }\]

OpenStudy (anonymous):

?

Parth (parthkohli):

Add the numerators...

hero (hero):

@ParthKohli, see was asking if \[(1 - \cos ^2 \theta) + (1-\cos^2\theta)\] is the deominator. No it is not.

OpenStudy (anonymous):

@Hero what do you mean??

OpenStudy (anonymous):

@ParthKohli oh, \[\frac{ 4 }{ 1-\cos ^2\theta } ?\]

Parth (parthkohli):

Yes! So we were checking if\[\dfrac{4}{1 - \cos^2 \theta} = 4\sec^2\theta\]Divide both sides by \(4\).

OpenStudy (anonymous):

no, its not sec, its csc

Parth (parthkohli):

I have a bad eyesight. Nvm. Can you divide both sides by \(4\)?

Parth (parthkohli):

For the easiness of life,\[\dfrac{4}{1 - \cos^2 \theta} = 4\times\dfrac{1}{1 - \cos^2\theta}\]

OpenStudy (anonymous):

I don't get it. How come that the \[4\csc^2\theta became 4 \times \frac{ 1 }{ 1-\cos^2\theta }?\]

hero (hero):

My only question is, @CHAsio_P, what process you have to perform in order to find the denominator?

Parth (parthkohli):

No, look you have:\[\frac{4}{1 -\cos^2\theta} = 4\csc ^2 \theta \]Look at the left hand side.

OpenStudy (anonymous):

@ParthKohli Oh, sorry. Then, will I already chnage the \[4 \csc^2 \theta \to 4(\frac{ 1 }{ \sin^2\theta } ?\]

Parth (parthkohli):

Yes :-)

Parth (parthkohli):

And another thing: Do you know how you could simplify \(1 - \cos^2\theta\)?

Parth (parthkohli):

There is a well known Pythagorean Identity associated with it.

OpenStudy (anonymous):

Yup

hero (hero):

@ParthKohli, I don't mean to interrupt, but when she asked: "So is the denominator \((1 - \cos ^2 \theta) + (1-\cos^2\theta)\)", you agreed with her.

OpenStudy (anonymous):

So, \[\frac{ 4 }{ \sin^2\theta } = 4 \frac{ 1 }{ \sin^2\theta } which is \frac{ 4 }{ \sin^2\theta } ?\]

Parth (parthkohli):

Sorry, I said "Yup" to the line I wrote just after :-P

Parth (parthkohli):

Yes @CHAsio_P!

OpenStudy (anonymous):

@Hero its my mistake. Don't worry about it. :)) I already got it.

Parth (parthkohli):

And how would you simplify\[\dfrac{4}{1-\cos^2\theta}\]

OpenStudy (anonymous):

Oh my gosh! @ParthKohli thanks for your help!! :))))

hero (hero):

What do you mean you already got it? I'm not convinced you have it.

Parth (parthkohli):

You're welcome (I'm not really a great teacher!)

OpenStudy (anonymous):

Well,its \[\frac{ 4 }{ \sin^2\theta } for 1 - \cos^2\theta is equals \to \sin^2\]

Parth (parthkohli):

Right there.

OpenStudy (anonymous):

@Hero it is just the denominator. I was just confuse, okay?

OpenStudy (anonymous):

@ParthKohli thanks again!!!

Parth (parthkohli):

No problem. :-D

hero (hero):

So @CHASsio_P, can you explain in your own words, the process of finding the denominator?

OpenStudy (anonymous):

Hmm. . . get the LCD. Heheh... I just dont know how to explain it by words, but i already got it. I swear. And I dont worry about the denominator, it is really my weakness. :))

hero (hero):

Here's a simple question: How do you find the LCD?

OpenStudy (anonymous):

Get the least common denominator or multiply the denominators .:))

hero (hero):

So if you had \[\frac{a}{b} + \frac{c}{d} = 5\] how do you find the LCD of that?

OpenStudy (anonymous):

Oh, \[\frac{ ad }{ bd }+\frac{ ac }{ bd } ?Am I correct?\]

hero (hero):

Yes, that is correct. Earlier you posted that for your original problem, the denominator should be \[(1 - \cos ^2 \theta) + (1-\cos^2\theta)\]. Can you correct your mistake here? What should that be instead?

OpenStudy (anonymous):

Oh yeah. . . the sign, I am sorry okay. I was just confused. :))

hero (hero):

Wait, actually, what you posted is not correct

hero (hero):

\[\frac{ad}{bd} + \frac{ac}{bd}\] is not correct. And you made more than just a sign mistake for \[(1 - \cos ^2 \theta) + (1-\cos^2\theta)\]

hero (hero):

First, for \[(1 - \cos ^2 \theta) + (1-\cos^2\theta)\], you should not be adding them together.

hero (hero):

And then, yes you're right, there is a sign mistake. Instead of adding, multiply so what we should have for the denominator is \[(1 - \cos ^2 \theta) (1+\cos^2\theta)\] and of course that is equal to \[1 - \cos^2\theta\]

OpenStudy (anonymous):

Okay. Thanks for explaining. :))

Parth (parthkohli):

@Hero: Nice work :-)

hero (hero):

Also for the other fraction, you want: \[\frac{ad}{bd} + \frac{bc}{bd}\]

hero (hero):

I think you might be able to figure things out from here, but for the future, hopefully, you'll try to understand how you might explain the process of finding an LCD or any math concept you've learned to others because only when you can explain it is when you have truly learned it.

OpenStudy (anonymous):

I agree with you.

hero (hero):

It's not necessarily enough to demonstrate computational ability. Being able to express the process verbally or written is the true measure of understanding.

hero (hero):

But anyway, I'm done lecturing. Good luck to you :D

OpenStudy (anonymous):

Yeah, thanks!! :))